OpenStudy (anonymous):

calculate the rotational kinetic energy of all molecules in one mole air at25 celsius

OpenStudy (anonymous):

Thr formula is T = ( 2 / 3k ) < K.E > T=25 , k= 1.38 x 10^-23 JK^-1 So 25 = ( 2 / 3 x 1.38 x 10^-23 ) <K.E> 25 = ( 2 / 4.14 x 10^-23 ) <K.E> 25 x 4.14 x 10^-23 = 2 x <K.E> 1.1025 x 10^-21 = 2 x <K.E> (1.1025 x 10^-21) / 2 = <K.E> 5.5125 x 10^-22 = <K.E>